OpenStudy (anonymous):

commodities x and y are perfect complements and the utility function is given by u =min(x,3y) price of x is 2 price of y is 1 income =140 find equilibrium level of consumption

OpenStudy (anonymous):

Although we could solve this by means of the lagrangean - in this case you're given everything you'll need to make this a simple multivariate within the question. Since we know that the goods are perfect complements and the Utility function is given by U=min(x,3y) we can infer that regardless of income - in the end your consumption bundle will be x=3y (meaning you take 3 units of good x for each unit of y) So x=3y. At the same time, we know that the budget constraint can be defined as 2x+y=140. Simple substitution yields this : 2(3y)+y=140 7y=140 y=20 x=3y x=60 Nice and simple, eh?

OpenStudy (anonymous):

i also thought the same but it was supposed to be a tough question and i thought answer could not be a simple linear equation. yeah but it makes perfect sense so thanks man.